OpenStudy (anonymous):

This problem studies such effects and how a Zener diode may be used to ameliorate the problem. Assume that VI=13.5V, RIN=1 kΩ, and vi=55.0mV. how do i workout the value of vo

OpenStudy (anonymous):

Hi, this worked for me - vo=vin/Rin *(RD*RL)/(RD+RL) so Try this values for both vo : 0.055/1000* (1*2000)/(2001) So for both RL 2k and 4k you should have ~ 0.000055

OpenStudy (anonymous):

how do i work out rd sorry

OpenStudy (nick67):

did you check here: http://openstudy.com/study#/updates/5062c622e4b0da5168bd6c57 it's quite similar

OpenStudy (anonymous):

i did but there wasnt a visible formula i could use

OpenStudy (anonymous):

RD should be 1, value taken from the graphic. Did my answers were correct?

OpenStudy (nick67):

noise value is: vo = vi*Rz/(Rz+Rin), where Rz is 1 Ohm

OpenStudy (anonymous):

0.000055 was right for 4k but not 2k

OpenStudy (nick67):

the noise output with zener is quite not dependent by the load

OpenStudy (nick67):

as far as the Rload is much higher than Zener resistance, load contribution to vo calculation is negligible

OpenStudy (anonymous):

hmmm it should be almost the same, try close values :0.000054, 0.000056 should be fine

OpenStudy (anonymous):

iv attached a print screen so you can see if i have done summit wrong

OpenStudy (nick67):

I was talking about values with Zener in

OpenStudy (nick67):

with RL 2K vo about 54.92 uV, with RL 4K vo about 54.93 uV

OpenStudy (anonymous):

oooohh this is the first part, where there isn't zener. The formula is a bit different here. Try these values: for 2k - 0.037 and for 4k 0.044 is it ok?

OpenStudy (anonymous):

yes you both are life savers i was stuck on this for the past 9 hours one last question is how do we work out the minimum rl

OpenStudy (nick67):

without Zener, vo is 36.67 mV for RL = 2k and 44 mV for RL = 4K

OpenStudy (anonymous):

The formula for RL is : Vin*RL/(RL+Rin)>=5V

OpenStudy (anonymous):

Try 580-590 value, ok?

OpenStudy (anonymous):

thank you for the help mate best respone and this has probably helped alot of others that have been stuck on this one

OpenStudy (anonymous):

so it's everything ok now?

OpenStudy (anonymous):

uhhhh how did you find V0 (capital V)

OpenStudy (anonymous):

Try this if you are asking about the first part - VO=VI*RL/(RL+RIN) If not give more info what you need.